# GOAL: A Generalist Combinatorial Optimization Agent Learner

### v1

@@ 4.1 Performance on the Training Tasks
BQ-NCO greedy 1.63% 1s 1.55% 9s 4.29% 26s 8.09% 1m
GOAL multi-task greedy 0.45% 5s 1.54% 30s 1.48% 3m 1.89% 1h
@@
GOAL multi-task greedy 2.85% 5s 2.80% 12s 3.63% 48s 6.63% 3m
JSSP10x10 JSSP15x15 JSSP20x15 JSSP20x20
COMPASS 4.70% 23m 8.00% 38m 10.40% 1h -
GOAL multi-task greedy 2.65% 6m 5.90% 18m 17.40% 32m 10.80% 59m
@@
4.2 Fine-Tuning to New Tasks
In this section, we want to evaluate the transfer capability of the trained GOAL backbone to unseen tasks. We consider four new CO problems: the Traveling Repairman Problem (TRP), the Open Vehicle Routing Problem (OVRP), the Split Delivery Vehicle Routing Problem (SDVRP) and the Maximum Independent Set (MIS) problem (detailed description in Appendix A). For each problem, our first attempt was to define new task-specific input and output adapters and train them either with or without back-propagation through the backbone. It turned out that in that case, the fine-tuning process is very unstable and has a strong dependency on the parameter initialization (i.e. the random seed). Next, we initialized the input adapters with the ones of the closest training problems: TSP for TRP, CVRP for OVRP and SDVRP and MVC for MIS. The output adapters were still trained from scratch. In this case the fine-tuning is very effective. In Fig 2, we see that the finetuned model reaches very good optimality gaps (at most $5\%$) much faster than a single-task GOAL model trained from scratch. Our few-shot fine-tuning process relies on only 64 instances and lasts less than 1 minute. We use the same loss and optimizer as for the training but with a smaller initial learning of 0.0001. These results assume that the whole model is fine-tuned, including the backbone. When the latter is frozen, we observe a slight degradation of the results with no significant speedup. From these experiments, we hypothesize that the input adapters are the difficult part of the transfer. To test this hypothesis, we considered one of the training tasks, discarded its input and/or output adapters and tried to re-learn them from scratch, with the trained backbone. This led to similar observations: re-learning the output adapter works, while re-learning the input adapter is very unstable. We conclude that transfer in GOAL is successful for problems that share the same instance structure as one of the training problems and thus can benefit from a pre-trained input adapter. Note that the new problems can have arbitrary objective and constraints (as is the case with the four tested new tasks). Appendix D further discusses input adapters and their impact on the latent representations.
Figure 2: Fine tuning GOAL to new tasks.
4.3 Ablation Study
Figure 3: Comparison of the model performance when using different variants of mixed-attention blocks for solving the Asymmetric Traveling Salesman problem
Figure 4: Comparison between single and multi-type architectures for the Job Shop Scheduling Problem
Mixed-attention blocks.
To evaluate the effectiveness of our mixed-attention blocks, we replace them by alternative modules which also aim at mixing the edge information with the attention scores, namely MatNet [34] and G2G [20]. More precisely we focus on the ATSP and train three versions of GOAL with the three different blocks. For MatNet, we keep the original hyper-parameters used for solving the ATSP. Figure 4 presents the models performance (optimality gap on a test set of 128 instances) at different points of the training. We see that using our mixed-attention blocks leads to a significant gain, on the performance at convergence (MatNet) or on the speed of convergence (G2G). Note that these ablations are performed on a smaller training dataset than in the previous experiments (100k instances), hence the wider opt gap.
Multi-type transformer.
The proposed multi-typed transformer architecture is meant to deal with tasks that are naturally represented by multipartite graphs with heterogeneous nodes or edges, typically scheduling problems where jobs and machine nodes play different roles. To evaluate its effect on the performance, we compare it to that obtained with the more direct alternative which ignores node heterogeneity: we simply concatenate the features of the different types of nodes (or edges) and use dummy values whenever a feature is undefined, so that the original single-type model can be used. We implemented this alternative for the Job Shop Scheduling Problem and present in Fig 4 a comparison between the training performance of the single-type and multi-type architectures. Clearly the multi-type architecture converges faster and to a better performance: after 200 epochs of training on 100K instances, it reaches a test optimality gap of 4.6% while the single-type model achieves only 8.9%.
5 Discussion and Conclusion
In this paper, we propose a generic neural CO model (GOAL) and demonstrate the feasibility of learning a shared backbone model to solve a variety CO problems, spanning routing, scheduling, packing and classic graph problems. We have shown the benefits of GOAL as (i) a single-task specialized model, providing state-of-the-art results on six CO problems; (ii) a multi-task model, able to solve several problems with the same backbone; and (iii) a pretrained model, to be efficiently fine-tuned for new problems. To the best of our knowledge, this is the first neural approach that handles such a variety of problems with a single backbone model. We believe this is a key step in the direction of a foundation model for CO, where the diversity of problem representations (as opposed to the uniformity of textual or visual data) is still a major hurdle.
Our experiments have exclusively been conducted with Imitation Learning, which assumes the availability of good-quality solutions for at least a certain size of instances. While this is not an intrinsic limitation of the GOAL model – the underlying MDPs being defined independently of the training process – we believe that Imitation Learning is an effective training paradigm for learning multi-task models. This is indeed demonstrated by previous works which relied on sequence-modeling training tasks for control [51, 47]. Leveraging solutions to classical NP-hard problems, provided by traditional solvers, which are based on decades of OR research, is particularly advantageous for pretraining a model. Such a model can later be adapted to more exotic problems with minimal supervision data, or even better, with Reinforcement Learning. Indeed using Reinforcement Learning to fine-tune GOAL to new problems, for which high-quality solvers may not be available, constitutes a promising direction for future work.
Appendix A Description of the considered CO problems
In this Section, we describe the CO problems that are solved by GOAL as well as the instance generation process. We organize the problems into three categories: Routing, Scheduling and Packing and Graph Problems. We also provide an overview Table of the problems and used solvers for each category.
A.1 Routing Problems
TSP TRP OP
Node features - - Price
Edge features Distance between nodes
Instance features - - Maximum tour length
Objective $min$(tour length) $min$(sum of waiting times) $max$(collected prices)
Oracle Concorde(LKH) LKH EA4OP
Action Select a node
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CVRP/SDCVRP CVRPTW OCVRP
Node features Demands Demands Time window starts Time window ends Service times Demands
Edge features Distance between nodes
Instance features Vehicle capacity
Objective Minimize total tour length
Oracle Concorde/HGS HGS LKH
Action Select a next node, subroute is done
[Uncaptioned image] [Uncaptioned image]
Table 3: Overview of the routing problems solved by GOAL
A.1.1 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) is a classic combinatorial optimization problem. The objective of the TSP is to find the shortest possible route that a salesman can take to visit a given set of cities exactly once and return to its starting point. In the path-TSP variant, the origin and destination of the salesman are distinct nodes (but may have the same coordinates, so that the TSP is a special case of the path-TSP). A TSP instance can be represented either by node coordinates or by a distance matrix. The majority of NCO models address TSP in the coordinate setting, yet the matrix setting holds greater significance for practical applications, as it applies beyond the Euclidean context, where the matrix can represent time, energy or cost, rather than distance, and need not satisfy the symmetry property of a distance matrix.
Data generation
For the Euclidian case, we generate datasets as described in [30]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed. For non-Euclidean problems (ATSP) we generate data as described in [34], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
A.1.2 Capacitated Vehicle Routing Problem
@@
The Capacitated Vehicle Routing Problem (CVRP) is another standard routing problem. The goal of CVRP is to optimize the delivery routes of a fleet of vehicles to a set of customers, while considering each vehicle’s capacity. The objective is to minimize the overall distance traveled by the vehicles, while respecting the capacity constraints and ensuring that all the customer demands are met.
@@
The matrix is sampled in the same way as for TSP, while node demands are randomly generated integers from the interval [1, 10]. Following [30], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
A.1.3 CVRP with Time Windows
The CVRP with Time Windows (CVRPTW) is an extension of the CVRP. In addition to demand constraints, each customer also has a specified time window during which they can be visited, along with the necessary service time required for serving that customer. The objective of the CVRPTW is to find routes for the vehicles that satisfy the capacity limits, deliver goods within the specified time windows, and minimize the total travel and waiting time or the number of vehicles used. In our model, we consider the standard version of the problem where total traveling and serving time should be minimized.
@@ A.1.3 CVRP with Time Windows
For the time windows and service times, we use a similar procedure to [17], where the travel times between nodes are sampled in the range [0, $\sqrt{2}\pm 0.1$]. We ensure the feasibility of the resulting instances by sampling time windows using a more refined procedure. Windows start time values are sampled from the range (0.01, 0.25). It’s important to note that windows start time do not directly impact problem feasibility, as vehicles can wait at the node until the service starts. On the other hand, the choice of the windows end time has an impact on feasibility. An instance is unfeasible if the window end time of a customer is less than the travel time from the depot plus the service time. To avoid this, windows end time are sampled from the interval [max(window start time, travel time from depot) + 2 $\times$ service time, max(window start time, travel time from depot) + 4 $\times$ service time].
A.1.4 Split Delivery VRP
The Split Delivery VRP (SDVRP) is similar to the CVRP with the difference that a customer can be visited by several vehicles, i.e. its demand is split between multiple vehicles.
@@ A.1.4 Split Delivery VRP
We use the same data as for the CVRP (Sec A.1.2).
A.1.5 Open Vehicle Routing Problem
This is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers, and they do not need to return to the depot at the end of their service.
@@ A.1.5 Open Vehicle Routing Problem
We use the same data as for the CVRP (Sec A.1.2).
A.1.6 Orienteering Problem
@@
Again we follow previous work [30]: node locations are sampled in the same way as for the TSP and CVRP and prizes are randomly sampled integers from the interval [1, 10]. The maximal distance constraint is set to 4.
A.1.7 Traveling Repairman Problem
In the travelling repairman problem (TRP), as in TSP, a set of locations is given as well as the time needed to travel between any pair of locations. Each location has a task to be performed by the repairman. The objective of the problem is to find the route that minimizes the sum of waiting times for all tasks.
@@ A.1.7 Traveling Repairman Problem
We use the same data as for the TSP (Sec A.1.1)
A.2 Packing and Graph Problems
KP MVC MIS
Node features Price and weight -
Edge features - Adjacency matrix
Instance features Knapsack capacity -
Objective $max$(collected prices) $max$(vertex coverage) $min$(independent set)
Oracle ORTools FastWVC
Action Select a node
Table 4: Overview of the packing and graph problems solved by GOAL
A.2.1 Knapsack Problem
The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set. Each item had a weight and a value, and the goal is to maximize the cumulated value of selected items, but without exceeding a given cumulated weight threshold, known as the knapsack’s (weight) capacity.
@@ A.2.1 Knapsack Problem
Following the same procedure as in [33], node weights and values are randomly sampled from the unit interval and knapsack capacity is set to 25 for all problem sizes.
A.2.2 Minimum (Weighted) Vertex Cover Problem
The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices. In the weighted version of this problem (MWVC), each vertex has an associated weight, and the goal is to select vertices with the minimum cumulated sum of weights while satisfying the vertex cover condition.
@@ A.2.2 Minimum (Weighted) Vertex Cover Problem
For this problem, we follow a similar procedure as in [27]. We generate random graphs by using Erdős–Rényi model with the critical edge probability $p$ sampled from [0.05, 0.15].
A.2.3 Maximum Independent Set
In the Maximum Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge.
Data generation.
We use the same data generation process as for the Minimum (Weighted) Vertex Cover Problem.
A.3 Scheduling Problems
UMSP JSSP
Node features Machine: Availability time Task: Execution time, Availability time Machine: Availability time
Edge features Job-machine: Execution times Task-task: precedence constraints, $ij=1$ if tasks $i$ and $j$ belong to the same job Task-machine: $ij=1$ if task $i$ must be executed on machine $j$
Objective Minimize a makespan
Oracle HiGHS ORTools
Action Select a machine for a random task Select a task
[Uncaptioned image] [Uncaptioned image]
Table 5: Overview of the scheduling problems solved by GOAL
A.3.1 Job Shop Scheduling Problem
The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem. We are given $N$ jobs and $M$ machines. Each job consists of a sequence of operations that need to be processed in a specific order (implying precedence constraints). Each operation has a given processing time. Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption). A solution is a schedule, defined by the start time of each operation such that the makespan, i.e. the maximum completion time of the operations, is minimized.
Data generation.
We follow the standard procedure for generating JSSP benchmark datasets: execution times are randomly sampled from $U(1,10)$.
A.3.2 Unrelated Machine Scheduling Problem
In the Unrelated Machine Scheduling problem (UMSP), the goal is to schedule $N$ jobs on $M$ machines such that the makespan is minimized. Each job $i$ has a processing time $p_{ij}$ on each machine $j$, which is arbitrary (hence the term unrelated). There are no precedence constraints and the schedules are non-preemptive. The solution to the problem is essentially a partition of the jobs on the machines. [21] shows that this problem is NP-hard.

### v2

@@ Appendix D Generalization to new instance distributions
BQ-NCO greedy 1.63% 1s 1.90% 9s 4.29% 26s 8.09% 1m
GOAL multi-task greedy 0.50% 5s 1.84% 53s 1.77% 3m 2.37% 28m
@@
GOAL multi-task greedy 3.06% 5s 3.35% 30s 3.96% 2m 7.63% 3m
@@
Appendix E Description of the considered CO problems
In this Section, we describe the CO problems (in alphabetical order) that are solved by GOAL as well as the instance generation process. GOAL solves problems from various categories: Routing, Scheduling, Packing, Graph, and Location Problems. For each problem, we provide all the details of how it is represented in our model, including a list of instance, node and edge features, the action chosen by the model, and the objective. We also describe how we generate a solution trajectory for problems where we use single cross-entropy loss for behavioral cloning.
E.1 Capacitated Vehicle Routing Problem and Split-Delivery Capacitated Vehicle Routing Problem
@@
The Capacitated Vehicle Routing Problem (CVRP) is a standard routing problem. The goal of CVRP is to optimize the delivery routes of a fleet of vehicles to a set of customers while considering each vehicle’s capacity. The objective is to minimize the overall distance traveled by the vehicles while respecting the capacity constraints and ensuring that all the customer demands are met. Split-Delivery CVRP (SDCVRP) is a variant of the CVRP that allows a single customer’s demand to be fulfilled by multiple vehicles. In our experiments we use Euclidean CVRP for training GOAL and SDCVRP for fine-tuning.
@@
Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10]. Following [kool_attention_2019], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
Instance features: Remaining vehicle capacity
Node features: Origin/destination tokens, demand
Edge features: Distance between nodes
Oracle: LKH
Objective: Minimize traveled distance
Action: Select a node, with direct edge or via-depot
Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
remaining capacity in descending order
E.2 Capacitated Vehicle Routing Problem with Time Windows
The Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) is an extension of the CVRP. In addition to demand constraints, each customer also has a specified time window during which they can be visited, along with the necessary service time required for serving that customer. The objective of the CVRPTW is to find routes for the vehicles that satisfy the capacity limits, deliver goods within the specified time windows, and minimize the total travel and waiting time or the number of vehicles used. In our experiments, we consider the standard version of the Euclidean problem where total traveling and serving time should be minimized and we use it for training GOAL.
@@ E.2 Capacitated Vehicle Routing Problem with Time Windows
For the time windows and service times, we use a similar procedure to [falkner_learning_2020], where the locations are sampled sampled from the unit square. We ensure the feasibility of the resulting instances by sampling time windows using a more refined procedure. Windows start time values are sampled from the range (0.01, 0.25). It’s important to note that start times do not directly impact problem feasibility, as vehicles can wait at the node until the service starts. On the other hand, the choice of the visiting end time has an impact on feasibility. An instance is unfeasible if the window end time of a customer is less than the travel time from the depot plus the service time. To avoid this, end times are sampled from the interval [max(window start time, travel time from depot) + 2 $\times$ service time, max(window start time, travel time from depot) + 4 $\times$ service time].
Instance features: Remaining vehicle capacity, current time
Node features: Origin/destination tokens, demand, start time window, end time window,
service time
Edge features: Distance between nodes
Oracle: HGS
Objective: Minimize traveled time
Action: Select a node, with direct edge or via-depot
Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
remaining capacity in descending order
E.3 Job Shop Scheduling Problem
The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem. We are given $N$ jobs and $M$ machines. Each job consists of a sequence of operations (they can called tasks) that need to be processed in a specific order (implying precedence constraints). Each operation has a given processing time. Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption). A solution is a schedule, defined by the start time of each operation such that a specific criterion, such as makespan, total completion time, or total tardiness is minimized.
@@ E.3 Job Shop Scheduling Problem
We follow the standard procedure for generating JSSP benchmark datasets: execution times are randomly sampled from $U(1,100)$.
Instance features: None
Node features: Job execution time, job availability time, machine availability time
Edge features: Precedence constraint, task-machine dependency
Oracle: ORTools
Objective: Minimize makespan
Action: Select a task
Trajectory: Tasks ordered by minimal finishing time
E.4 Knapsack Problem
The Knapsack Problem (KP) is a well-known combinatorial optimization problem that involves selecting a subset of items from a given set. Each item had a weight and a value, and the goal is to maximize the cumulated value of selected items, without exceeding a given cumulated weight threshold, known as the knapsack’s (weight) capacity.
@@ E.4 Knapsack Problem
Following the same procedure as in [kwon_pomo_2020], node weights and values are randomly sampled from the unit interval, and knapsack capacity is set to 25.
Instance features: Remaining knapsack capacity
Node features: Value, weight
Edge features: None
Oracle: ORTools
Objective: Maximizing packed values
Action: Select an item
Trajectory: There is no order
E.5 Maximal Independent Set Problem
In the Maximal Independent Set (MIS), we are given an undirected graph $G=(V,E)$ and the goal is to find a subset of nodes $S\subset V$ of maximal size $|S|$ such that each pair of nodes in $S$ is not connected by an edge. Simple speaking, a Maximal Independent Set (MIS) in a graph is a subset of vertices that is independent (no two vertices in the set are adjacent) and is not a subset of any larger independent set. In our experiments, we fine-tune GOAL to MIS problem.
Data generation.
For this problem, we follow a similar procedure as in [khalil_learning_2017]. We generate random graphs by using Erdős–Rényi model with the critical edge probability $p$ sampled from [0.05, 0.15].
Instance features: None
Node features: None
Edge features: Adjacency value (0-1)
Oracle: FastWVC
Objective: Maximize number of selected nodes
Action: Select an node
Trajectory: There is no order
E.6 Maximum Coverage Location Problem
In the Maximum Coverage Location Problem (MCLP), we are given a set of locations with associated weights and several facilities with associated radius of coverage. The goal is to find the optimal placement of the facilities, on a subset of the locations, to maximize the number of covered locations. There is also a variant in which nodes have associated weights; in this case, the objective is to maximize the sum of the weights of the covered locations. In our experiments, we fine-tune our multitask model to MCLP.
Data generation.
Node locations are randomly sampled from the unit square, the number of facilities is fixed to 10 and their radius of coverage to 0.1.
Instance features: Number of facilities, radius of coverage
Node features: None
Edge features: Distance between nodes
Oracle: PuLP [^2] https://github.com/coin-or/pulp
Objective: Maximize the number of covered locations
Action: Select an facility node
Trajectory: There is no order
E.7 Minimum Vertex Cover Problem
The goal of the Minimum Vertex Cover Problem (MVC) is to find the smallest set of vertices in a graph, such that each edge is incident to at least one of these vertices. In the weighted version of this problem (MWVC), each vertex has an associated weight, and the goal is to select vertices with the minimum cumulated sum of weights while satisfying the vertex cover condition. Data generation and problem features are the same as in MIS. We use MVC to train our model.
E.8 Open Shop Scheduling Problem
The Open Shop Scheduling Problem (OSSP) is a scheduling problem in which a set of jobs (with a given set of operations) must be processed on a set of machines, but unlike job shop scheduling, there is no predefined order. Each operation must be processed, with no overlapping between operations and machines at the same time unit. The goal is typically to minimize the makespan, or the total time required to complete all jobs. Data generation and problem features are the same as in JSSP, except OSSP does not have precedence constraints.
E.9 Open Vehicle Routing Problem
Open Vehicle Routing Problem (OCVRP) is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers. In contrast with other CVRP variants, vehicles do not need to return to the depot at the end of their service. This makes OCVRP similar to the classical CVRP problem, with a different objective function. In our experiments, we handle the Euclidean version of the OCVRP and fine-tune GOAL to solve it.
Data generation.
Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10]. As in CVRP, the vehicle capacity is set to 50 for a problem size of 100.
Instance features: Remaining vehicle capacity
Node features: Origin/destination tokens, demand
Edge features: Distance between nodes
Oracle: LKH
Objective: Minimize traveled distance
Action: Select a node, with direct edge or via-depot
Trajectory: Nodes ordered by a solution subtour. Subtours are ordered by their final
remaining capacity in descending order
E.10 Orienteering Problem
@@
We generated data following previous work [kool_attention_2019] and  [xin_multi-decoder_2021]: node locations are uniformly sampled from the unit square and prizes are proportional to the distance to the depot. We followed the same procedure line in  [kool_attention_2019], this variant of the problem is called $OP_{distance}$. The maximal tour length constraint is set to 4. We use Euclidean OP for training GOAL.
Instance features: Remaining tour length
Node features: Origin/destination tokens, prize
Edge features: Distance between nodes
Oracle: A4OP
Objective: Maximize collected prizes
Action: Select a node (including coming back to the depot)
Trajectory: Ordered nodes by a solution tour
E.11 Price Collection Traveling Salesman Problem
The Prize-Collecting Traveling Salesman Problem (PCTSP) is a variant of the classic Traveling Salesman Problem. In the PCTSP, each node has an associated prize and penalty, and the goal is to minimize the traveled distance while satisfying a minimum total prize constraint. Unlike the traditional TSP, where all cities must be visited, the PCTSP allows for skipping some nodes; however, unvisited nodes incur a penalty, creating a trade-off between collecting prizes and minimizing travel distance. The final objective is to minimize the sum of the traveled distance and the sum of the penalties for unvisited nodes.
@@ E.11 Price Collection Traveling Salesman Problem
We generated data following previous work [xin_multi-decoder_2021]: node locations are uniformly sampled from the unit square, while prizes are randomly sampled from $U[0,num-nodes]$ and penalties are sampled from $U[0,num-nodes]$. Both features are normalized by multiplying by $4/num-nodes$ and $12/num-nodes$, respectively.
Instance features: Remaining prize constraint
Node features: Origin/destination tokens, prize, penalty
Edge features: Distance between nodes
Oracle: ORTools
Objective: Minimize the sum of traveled distances and penalties of univisited nodes
Action: Select a node (including coming back to the depot)
Trajectory: Ordered nodes by a solution tour
E.12 Sequential Ordering Problem
The Sequential Ordering Problem (SOP) is a combinatorial optimization problem where a set of tasks or nodes must be visited in a specific order while minimizing the total travel cost or distance. This problem can be modeled as the Traveling Salesman Problem, with precedence constraints. The objective is to find the shortest possible route that satisfies these precedence conditions. In our experiments, we fine-tune our multitask model to SOP.
@@ E.12 Sequential Ordering Problem
The values of the cost matrix are integers randomly generated from a uniform distribution $U[1,1000]$ with zero costs between the same task. To fulfill feasibility constraints, we randomly order the tasks, and for the $i$-th task, we randomly add $k$ precedent jobs, where $k$ is sampled from $U[0,\lfloor i/2\rfloor]$.
Instance features: None
Node features: Origin token
Edge features: Distance between nodes, precedence constraint
Oracle: LKH
Objective: Minimize cost
Action: Select a node
Trajectory: Ordered nodes by a solution tour
E.13 Traveling Repairman Problem
The Traveling Repairman Problem (TRP) is a combinatorial optimization problem where a repairman must determine the most efficient route to visit a set of locations, performing repairs or services at each site. In the Traveling Repairman Problem, a set of locations is given with the time needed to travel between any pair of locations. Each location has a task to be performed by the repairman. The objective of the problem is to find the route that minimizes the sum of the delays for reaching each point. In our experiments, we fine-tune GOAL to Euclidean TRP.
@@ E.13 Traveling Repairman Problem
Node locations are uniformly sampled from the unit square.
Instance features: None
Node features: Origin token
Edge features: Distance between nodes
Oracle: LKH
Objective: Minimize waiting time of all nodes
Action: Select a node
Trajectory: Ordered nodes by a solution tour
E.14 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) is a classic combinatorial optimization problem. The objective of the TSP is to find the shortest possible route that a salesman can take to visit a given set of cities exactly once and return to its starting point. In the path-TSP variant, the origin and destination of the salesman are distinct nodes (but may have the same coordinates, so that the TSP is a special case of the path-TSP). A TSP instance can be represented either by node coordinates or by a distance matrix. The majority of NCO models address TSP in the coordinate setting, yet the matrix setting holds greater significance for practical applications, as it applies beyond the Euclidean context, where the matrix can represent time, energy, or cost, rather than distance, and need not satisfy the symmetry property of a distance matrix. In our work, we use the Asymmetric TSP (where the distances satisfy the triangular inequality property) to train our multi-task model."
Data generation
For the Euclidean case, we generate datasets as described in [kool_attention_2019]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed. For non-Euclidean problems (ATSP) we generate data as described in [kwon_matrix_2022], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
Instance features: None
Node features: Origin/destination tokens
Edge features: Distance between nodes
Oracle: Concorde (for Euclidean) and LKH (for non-Euclidean)
Objective: Minimize tour length
Action: Select a node
Trajectory: Ordered nodes by a solution tour
E.15 Unrelated Machine Scheduling Problem
The Unrelated Machine Scheduling Problem (UMSP) is an optimization problem in operations research where a set of jobs must be assigned to a set of machines, each with different processing times for each job, to minimize a specific criterion, such as makespan, total completion time, or total tardiness. Unlike related machine scheduling, where machines have similar characteristics and processing times are proportional, in UMSP, the performance and efficiency of machines vary, making it a more complex and challenging problem to solve. [horowitz_exact_1976] shows that this problem is NP-hard. We use UMSP for multitask training of our GOAL model.